• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:03:01 on April 07, 2008

    Entry number 230558

    keyword=day shift summary

    Day shift summary:

    Shift workers: J. P. Chen (SL), X. Zhan (TO), X. Qian (3rd).

    8:00 Recovering from previous BigBite detector HVs trip. 8:15-9:00 taking data at kinematics A. HV tripped 2 more times during this period. 9:00 MCC working on beam. Moved to Empty. 9:10 Asked to have controlled access, need to wait at least 30 minutes. 9:30-9:45 changed to kinematics B: L_HRS to 16.5 degree, 1194 MeV. 10:30 full survey done. Hall in controlled access. People going to the hall: Bryan (for helicity), Doug (for Helium flow), Alex (for DAQ), Brad to get a box out with radcon clearance). 16:00 still no beam

    Run 2491/2492 Kinematics A. Both had HV trip at the end.